OpenStudy (anonymous):

The mass m(t) remaining after t years from 90-gram sample of a radioactive element is given by m(t)=90e^-0.3t. When will the mass remaining equal 9 grams? Round your answer to the nearest hundredth of a year.

OpenStudy (anonymous):

a. 2.30 years b. 7.68 years c. 0.12 years d. 3.33 years e. 0.13 years

OpenStudy (anonymous):

I substituted the t for each one of the answers but i dont get 9 with any of them

OpenStudy (slaaibak):

9 = 90e^(-0.3t) 1/10 = e^(-0,3t) ln (1/10) = -0.3t t = ln(10) / 0.3 t= 7.67528364 therefore t = 7.68

OpenStudy (anonymous):

Thanks mate